# Anomalous diffusion in convergence to effective ergodicity

### v3

@@ Abstract
The nature of diffusion is usually studied for particles or dynamics generating trajectories over time. Similar in principle, these studies can be executed in tracking how a given function of observable properties evolve over time akin to a system’s particle motion, so-called functional-diffusion. This is not the same as systems’ own trajectories but can be considered as a meta-trajectory. Following this idea, we measure how an approach to ergodicity evolves over time for the observed magnetization of a full Ising model with an external field. We compute the functional’s diffusive behavior depending on a range of temperatures via Metropolis and Glauber single-spin-flip dynamics. System’s ensemble-averaged dynamics are computed using expressions from the exact solution. Power-laws on the approach to ergodicity provide the classification of anomalies in the functional-diffusion, demonstrating non-linear super-diffusive behavior over different temperature ranges.
@@
ergodicity, power-laws, Ising models, lattice dynamics, monte carlo, functional-diffusion
@@ 1 Introduction
Brownian motion is probably one of the landmark concepts in statistical physics attracted Einstein’s interest early on [1]. Its importance in formulating statistical mechanics recently reviewed within its 250th anniverssary publications [2]. Major observable in tracing Brownian motion appears as how accumulated displacement curve behaves over time, i.e., linear relationship without intercept corresponding to normal diffusion. Along these lines, dynamics of cooperation among assembly of independent units has been studied in this context [3], such as model of magnetic material [4, 5, 6] and the state of a neuron [7, 8]. Measuring the meta-diffusion in this discrete cases are our core premise in an interesting setting: If the displacement curve of a functional of a function of trajectories shows a power-law scaling over time other than the exponent $1.0$, then the behaviour is called anomalous diffusion [9]. This setting not only yields to mathematically challenging consequences, provides insights into characteristics of the Ising model. Important to note that, described scenario does not apply to trajectories, rather a functional of trajectories. This distinction is the core driver in our work that a functional on the function that explains convergence to ergodicity for total magnetisation is the diffusing, rather than the discrete units of the Ising model.
@@
The formulation of the Ising model is reviewed in Sec. 2. In Sec. 3, the concept of effective ergodicity is explained. Measuring convergence to ergodicity is shown in Sec. 4; this corresponds to the functional-diffusion concept we introduced. Computation of power-law will be shown in Sec. 5. The conclusion is given in Sec. 6.
@@ 2 Ising-Lenz System
A lattice with $N$ sites, labelled as $\{s_{i}\}_{i=1}^{N}$ which can take two values, such as $\{1,-1\}$. These values imply spin up or down in the Ising Model [4, 5, 6] or an activation in neuronal systems [7, 8]. The total energy, the Hamiltonian of the system, can be written with two interactions, one due to nearest-neighbors (NN) and one due to an external field, with coefficients $J$ and $H$ respectively,
@@
Figure 1: Effective ergodic convergence in Metropolis and Glauber dynamics for $N=512$ with varying temperature and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
@@ 4 Ergodicity convergence as functional-diffusion
Figure 2: Effective ergodic convergence in Metropolis and Glauber dynamics for $N=1024$ with varying temperature and fixed external field $H=\{1.0\}$ at (a) and (b) respectively.
Effective ergodic convergence, $\Omega_{G}(t)$ for a given observable $g$, can be constructed based on the fact that identical parts of cooperating units have identical characteristics in thermodynamic equilibrium [11]. This is realised by the Thirumalai-Mountain (TM) $G$-fluctuating metric [11, 12], at a given time $t_{k}$
@@
Figure 3: Power-law exponents in effective ergodic convergence in Metropolis and Glauber dynamics for $N=512$ with varying temperature ($\beta$) and fixed external field $H=\{1.0\}$ at (a) and (b) respectively. Sharp increase at around $\beta=1.4$.
@@
$\Omega_{G}^{{}^{\prime}}=\frac{\Omega_{G}(t)}{\Omega_{G}(0)}\to\frac{1}{tD_{G}}$ (7)
where diffusion coefficient expressed as $D_{G}$. A similar measure of ergodicity is used in simple liquids [11, 22], and earthquake fault networks [23, 24].
@@
where $M_{T}(N,\beta,h)$ and $M_{E}(N,\beta,h)$ correspond to time-averaged and ensemble-averaged total magnetization. Note that exact expression for $M_{E}$ is used [10]. Hence, in computing $\Omega_{M}(t_{k},N,\beta,h)$, we generate Metropolis and Glauber single-spin-flip dynamics for $N=512$ cooperating units, a range of Boltzmann factor values $\beta=[0.5,1.5]$ with a spacing of $0.025$ and non-zero external field values $H=1.0$ with setting short-range interaction strength to $J=1.0$ [25, 26]. The time evolution of inverse rate, $\Omega_{G}^{-1}$, is given in Figure 1a and Figure 1b for $N=512$. After an initial period, inverse rate value decreases rapidly. At higher temperatures, inverse rate behaves differently. Similar results are produced for $N=1024$ on Figure 2a and Figure 2b.
Figure 4: Power-law exponents in effective ergodic convergence in Metropolis and Glauber dynamics for $N=1024$ with varying temperature ($\beta$) and fixed external field $H=\{1.0\}$ at (a) and (b) respectively. Sharp increase at around $\beta=1.4$.
5 Anomaly in the functional-diffusion
Inverse rate $\Omega_{G}^{{}^{\prime}-1}$ in the effective ergodic measure scales with Monte Carlo time in a power-law fashion,
$\Omega_{G}^{{}^{\prime}-1}\sim t^{-\alpha}$
where $\alpha$ is the exponent and $t$ is the Monte Carlo time. While power laws in complex systems and computing them on a given empirical data are studied in depth [27, 28]. We compute power-law exponents of evolution of inverse rate over Monte Carlo time. Identification where to start to compute $\alpha$ is achived by Kolmogorov-Smirnov test [29]. Temperature dependence of $\alpha$ is reported in Figure 3a and Figure 3b for $N=512$, where $\alpha>1.0$ is observed leading to superdiffusive behaviour. We also observe that there is a sharp increase for $\alpha$ at around $\beta=1.4$, along with a stable value in the range $\beta=(0.8,1.4)$. Due to different relaxation regimes, results deviate for $N=1024$, as seen in Figure 4a and Figure 4b, a parabola-like behaviour for superdussion strengths is observed over a temperature dependence.
@@ 6 Conclusion
A new concept of functional-diffusion is introduced via a canonical example on Ising model i.e., approach to ergodicity for total magnetisation. Superdiffusive regimes are identified based on different temperature ranges, via computation of power-laws of the inverse fluctuating metric for ergodicity. The results demonstrate size dependence on the relaxation dynamics and their anomalous diffusion for the functional cases. This concept would be a pedagogical test-bed for the extentions of diffusive behaviour beyond particle-trajectories.
Supplement
We provide two notebooks for reproducing the results, data_analysis.ipynb and data_generate.ipynb. Dataset is available on Zenodo [30]

### v4

@@ Abstract
The nature of diffusion is usually studied for particles or dynamics generating trajectories over time. Similar in principle, these studies can be executed in tracking how a given function of observable properties evolve over time akin to a system’s particle motion, so-called functional-diffusion. This is not the same as systems’ own trajectories but can be considered as a meta-trajectory. Following this idea, we measure how an approach to ergodicity evolves over time for the observed magnetization of a full Ising model with an external field. We compute the functional’s diffusive behavior depending on a range of temperatures via Metropolis and Glauber single-spin-flip dynamics. System’s ensemble-averaged dynamics are computed using expressions from the exact solution. Power-laws on the approach to ergodicity provide the classification of anomalies in the functional-diffusion, demonstrating non-linear anomalous behavior over different temperature and field ranges.
@@
ergodicity, power-laws, Ising models, lattice dynamics, monte carlo, functional-diffusion, Glauber dynamics, Metropolis dynamics
@@ 1 Introduction
Brownian motion is probably one of the landmark concepts in statistical physics attracted Einstein’s interest early on [1]. Its importance in formulating statistical mechanics recently reviewed within its 250th anniversary publications [2]. Major observable in tracing Brownian motion appears as how accumulated displacement curve behaves over time, i.e., linear relationship without intercept corresponding to normal diffusion. If the displacement curve of trajectories shows a power-law scaling over time other than the linear relationship, then the behaviour is called anomalous diffusion [3, 4, 5]. There is a recent surge in interest in using machine learning techniques in the analysis of anomalous diffusion data [6, 7, 8, 9], for example, the characterisation using transformer architectures is quite novel given the larger interests in attention mechanism in machine learning [10].
Along these lines, the dynamics of cooperation among assembly of independent units has been studied in this context [11], such as model of magnetic material [12, 13, 14] and the state of a neuron [15, 16]. Measuring the diffusion behaviour in this discrete cases would not only yields to mathematically challenging consequences, provides insights into characteristics of the Ising model. In our described scenario does not apply to trajectories, rather a functional of trajectories. This distinction is the core driver in our work that a functional on the function that explains convergence to ergodicity for total magnetisation is the diffusing, rather than the trajectory of the discrete units of the Ising model.
@@
The formulation of the Ising model is reviewed in Sec. 2. In Sec. 3, the concept of effective ergodicity is explained. Measuring convergence to ergodicity is shown in Sec. 4; this corresponds to the functional-diffusion concept we introduced. Computation of power-law will be shown in Sec. 5 and our experimental results in Sec 6. The conclusion is given in Sec. 7.
@@ 2 Ising-Lenz System
A lattice with $N$ sites, labeled as $\{s_{i}\}_{i=1}^{N}$ which can take two values, such as $\{1,-1\}$. These values imply spin up or down in the Ising Model [12, 13, 14] or an activation in neuronal systems [15, 16]. The total energy, the Hamiltonian of the system, can be written with two interactions, one due to nearest-neighbors (NN) and one due to an external field, with coefficients $J$ and $H$ respectively,
@@
Figure 1: Diagnostic plots of $40$ different runs for $N=1024$ with Glauber dynamics for the evolution of $K(t)$ log-log regression (a) A field at $H=1.0$ and inverse temperature $\beta=1.16$ (b) A field at $H=1.4$ and inverse temperature $\beta=1.0$. We identify optimal time starting fit with minimal Kolmogorov-Simirnov statistic on a grid-search.
Figure 2: Uncertainty quantified two power-laws’s $\alpha$ exponents over temperature ranges (a) the time-evolution of $K(t)$, via log-log regressions. (b) the distribution of $\Gamma(t)$ via analytical expressions from Newman et. al.. Here we plot absolute value as $\alpha$. In the fit, functional form is negative $\alpha$.
@@ 4 Ergodicity convergence as functional-diffusion
Effective ergodic convergence, $\Omega_{G}(t)$ for a given observable $g$, can be constructed based on the fact that identical parts of cooperating units have identical characteristics in thermodynamic equilibrium [18]. This is realized by the Thirumalai-Mountain (TM) $G$-fluctuating metric [18, 19]. Applications of ergodicity with TM metric is quite widely used also with recent interests, biomolecular simulations [29], physical chemistry and machine learning landscapes [30], seismology [31, 32, 33, 34, 35, 36], neuromorphic computing [37, 38] and artificial spin ice [39, 40].
TM metric at a given time $t_{k}$ reads,
@@
$\Gamma_{G}(t)=\frac{\Omega_{G}(t)}{\Omega_{G}(0)}\to\frac{1}{tD_{G}},$ (7)
where diffusion coefficient expressed as $D_{G}$. A similar measure of ergodicity is used in simple liquids [18, 41], and earthquake fault networks [42, 43]. The inverse rate of ergodic convergence can also be defined by $K_{G}(t)$,
$K_{G}(t)=\frac{\Omega_{G}(0)}{\Omega_{G}(t)}\to t\cdot D_{G}.$ (8)
@@
where $M_{T}(N,\beta,h)$ and $M_{E}(N,\beta,h)$ correspond to time-averaged and ensemble-averaged total magnetization. Since, ensemble averaged is computed analytically, this makes our approach using a modified TM-metric, leading to more accurate assessment of the ensemble averaging. Note that exact expression for $M_{E}$ is used [17].
5 Functional-diffusion: Two power laws
We would like to investigate the behaviour of the functional-diffusion in two power laws. The first-one is similar to the one investigated for displacement vector from the literature, but over ergodicity convergence function $K$ we defined above,
$K_{G}(t)\to C\cdot t^{\alpha}.$ (11)
whereby $C$ is the generalized diffusion coefficient, $t$ is the Monte Carlo time and for the positive $\alpha$ exponent. We call this time power-laws. The second type of power-law we seek is over the distribution of $\Gamma(t)$,
$P(\Gamma_{G}(t))\to\Gamma_{G}(t)^{-\alpha}.$ (12)
This is identified as distributional power-laws, whereby this is related to Lèvy flights or jump distributions [4], we consider cumulative jumps of approach to ergodicity here.
We will explicitly state which power-laws we are working on, not to confuse $\alpha$-s. The power laws in complex systems and computing them on a given empirical data are studied in depth using techniques pioneered by Newman et. al. [44, 45]. We did an extensive bias corrected bootstrapping for determining uncertainties on scaling exponents, diffusion coefficient, fit diagnostics, KS-distances (Kolmogorov-Simirnov Statistic) and adjusted R-squared, and autocorrelation times. KS-distance was critical in determining where to start to perform log-log regression depending on the parameters, this is done in an automated way with finding minimum KS-statistic over a search grid of time values. Visual inspection diagnostic is done for every parameters and diagnostic visualisations are developed and automated.
We also computed autocorrelations times using time self-correlations of averaged magnetisation, via location of the plateau of the correlation times $C(t)$ of averaged-magnetisations,
$C(t_{k})=\frac{1}{t_{k}}\sum_{i=0}^{k}M(t_{0})M(t_{i})$
. We identify that relaxation changes in the regions where anomalous convergence appears.
A Finite Size-Scaling (FSS) analysis could show a universality of the results depending on the number of discrete units and temperatures for the Ising model. We made the Following FSS assumptions, for temperature and size dependence
$u(\beta)=(\beta-\beta_{c})N^{a}$
scaling function $f(u)=A+Bu^{c}$, then power-law exponents $\alpha$ with the critical temperature $\beta_{c}$,
$\alpha(T,N)=N^{b}f(u)$
We identified the FSS exponents $a$, $b$, $c$ and coefficients $A$, $B$ with nonlinear optimization on the power-laws over temperatures for time power-laws.
Figure 3: Uncertainty quantified two power-laws’s $\alpha$ exponents over field ranges at $\beta=1.0$ (a) the time-evolution of $K(t)$, via log-log regressions. (b) the distribution of $\Gamma(t)$ via analytical expressions from Newman et. al.. Here we plot absolute value as $\alpha$. In the fit, functional form is negative $\alpha$.
Figure 4: Uncertainty quantified power-laws of the time-evolution of $K(t)$, via log-log regressions, diffusion coefficients (a) Over temperature ranges. (b) Over field ranges.
6 Experimental Results
We run extensive Metropolis and Glauber dynamics for the full Ising model and measure the rate of ergodic convergence and its inverse, $\Gamma(t)$ and its inverse $K(t)$. Here we present the result with our findings.
An of evolution of $K(t)$ is shown for 40 independent runs in Figure 1a and 1b for different temperature and field conditions, with it’s diagnostic of where we start fitting via optimal KS distance analysis. These plots served as a visualisal diagnostics in ensuring each temperature, size or field cases are inspected that the fitted log-log regression represents indeed a power-law region within the simulated MC time, whereby we record $\Gamma(t)$ are acceptance times.
Figure 5: Uncertainty quantified autocorrelation times over $200$ runs. (a) A diagnostic plot for identification of autocorrelation time. (b) Autocorrelation times over temperature ranges and three different sizes.
ikBT alphaMean alphaLower alphaUpper adjR2Mean adjR2Lower adjR2Upper
$0.50$ $0.0179$ $0.0145$ $0.0212$ $0.7133$ $0.6264$ $0.7830$
$0.56$ $0.0416$ $0.0368$ $0.0448$ $0.8870$ $0.8179$ $0.9186$
$0.62$ $0.0872$ $0.0818$ $0.0923$ $0.9710$ $0.9617$ $0.9769$
$0.68$ $0.1784$ $0.1726$ $0.1852$ $0.9815$ $0.9775$ $0.9847$
$0.74$ $0.3538$ $0.3462$ $0.3608$ $0.9897$ $0.9879$ $0.9913$
$0.80$ $0.6524$ $0.6407$ $0.6617$ $0.9939$ $0.9934$ $0.9946$
$0.86$ $1.1030$ $1.0812$ $1.1277$ $0.9969$ $0.9964$ $0.9972$
$0.92$ $1.4959$ $1.4892$ $1.5030$ $0.9995$ $0.9994$ $0.9995$
$0.98$ $1.7725$ $1.7665$ $1.7784$ $0.9999$ $0.9999$ $0.9999$
$1.04$ $1.8869$ $1.8780$ $1.8947$ $0.9998$ $0.9997$ $0.9998$
$1.10$ $1.8857$ $1.8754$ $1.8942$ $0.9995$ $0.9994$ $0.9996$
$1.16$ $1.8223$ $1.8110$ $1.8334$ $0.9992$ $0.9990$ $0.9993$
$1.22$ $1.7088$ $1.6951$ $1.7257$ $0.9990$ $0.9986$ $0.9991$
$1.28$ $1.5846$ $1.5678$ $1.6033$ $0.9983$ $0.9979$ $0.9986$
$1.34$ $1.4132$ $1.3864$ $1.4342$ $0.9966$ $0.9942$ $0.9975$
$1.40$ $1.1320$ $1.0944$ $1.1654$ $0.9876$ $0.9832$ $0.9907$
Table 1: Diagnostics summary for the power-law fit for the Glauber dynamics for log-log regression.
Power-law exponents alphaalpha for different sizes $N=512,1024,1536$ are identified for inverse temperature ranges. The log-log regression results are shown in Figure 2a for time power-laws, i.e., $K(t)$ observations over time. This is the functional-diffusion power-laws aligned with the conventional displacement power-laws. We observe anomalous behaviour super-diffusive and sub-diffusive regions, and normal diffusion region being in between. power-law has a different nature as shown in Figure 2b, it doesn’t necessarily represents diffusion behaviour, rather how Monte Carlo dynamics evolve. From simulations perspective this captures the dynamical behaviour due to Glauber spin-flip-dynamics, validating that anomalies present in the dynamics as well. More sophisticated cluster-flips maybe required for slow dynamics regions, such as the Swendsen-Wang [46] dynamics. Uncertainty estimates are achived using bias corrected bootstrapping for reliable confidence interval identification [47, 48]. We demonstrate a set of uncertainty quantification for Glauber dynamics $K(t)$ power-laws log-log plots in Table 1. The similar observations have been made for exponents $\alpha$ the field variations in Figures 3a and 3b. Generalized diffusion coefficients over time validates the same observations, 4a and 4b. Identification of the autocorrelation times from time correlations of average magnetisation. This indicates the different relaxation times matching with diffusion regimes we observed, in Figures 5a and 5b.
df type_power_law N D pvalue parameter
1 Distribution Temperature Range $512$ $0.1875$ $0.9522$ alpha scale
2 Distribution Temperature Range $1024$ $0.1875$ $0.9522$ alpha scale
3 Distribution Temperature Range $1536$ $0.1250$ $0.9998$ alpha scale
4 Distribution Field Range $512$ $0.2500$ $0.7164$ alpha scale
5 Distribution Field Range $1024$ $0.0625$ $1.0000$ alpha scale
6 Distribution Field Range $1536$ $0.1250$ $0.9998$ alpha scale
7 Time Temperature Range $512$ $0.0625$ $1.0000$ alpha scale
8 Time Temperature Range $1024$ $0.0625$ $1.0000$ alpha scale
9 Timer Temperature Range $1536$ $0.1250$ $0.9998$ alpha scale
10 Time Temperature Range $512$ $0.1250$ $0.9998$ C diffusion
11 Time Temperature Range $1024$ $0.1250$ $0.9998$ C diffusion
12 Timer Temperature Range $1536$ $0.1250$ $0.9998$ C diffusion
13 Time Field Range $512$ $0.0625$ $1.0000$ alpha scale
14 Time Field Range $1024$ $0.1875$ $0.9522$ alpha scale
15 Time Field Range $1536$ $0.1250$ $0.9998$ alpha scale
16 Time Field Range $512$ $0.0625$ $1.0000$ C diffusion
17 Time Field Range $1024$ $0.1250$ $0.9998$ C diffusion
18 Time Field Range $1536$ $0.1875$ $0.9522$ C diffusion
Table 2: Diagnostic summary for comparing power-laws on different ranges, with Glauber and Metropolis dynamics via two-sided KS test. Time means $K(t)$ power-laws and distribution means $P(\Gamma(t))$ power-laws.
Finite-size scaling parameters are also computed with nonlinear optimization. We found that $N=1024,1536$ behaves universally. It is also establised that results are statistically significant for both Metropolis and Glauber dynamics 2.
@@ 7 Conclusion
A new concept of functional-diffusion is introduced via a canonical example on Ising model i.e., approach to ergodicity for total magnetization. Superdiffusive regimes are identified based on different temperature and field ranges, via computation of power-laws of the forward and inverse fluctuating metric for ergodicity. Quite comprehensive results demonstrate for the first time anomalous convergences to ergodicity quantitatively with power-laws. This concept would be a pedagogical test-bed for the extensions of diffusive behaviour beyond particle-trajectories, so called the functional diffusion for full Ising model and increase the awareness on possible anomalous convergence of metric functionals.
Data Availability
Main diagnostic tables, all data generation and analysis R notebooks are available in public domain at a Zenodo repository [49].
Ackowlegements
We are grateful to Y.Süzen for her kind support of this work. Author would like to express the gratitude to referees for constructive recommendations, making the computational work more robust and pointing out relevant literature.